# Predicting small molecules solubilities on endpoint devices using deep ensemble neural networks

### v1

@@ 1 Introduction
Solubility prediction is critical and has driven the development of several methods, including first principles[10, 11], semi-empirical equations[12, 13, 14], molecular dynamics (MD) methods[15, 16, 17, 18], quantum computations[19], and quantitative structure-property relationship (QSPR)[20, 21, 22, 23] methods. Despite significant progress, the development of accurate and reliable models for solubility remains a major concern.[24]
To address the persistent issues of systematic bias and non-reproducibility in aqueous solubility datasets, Llinàs et al.[25, 26] introduced two solubility challenges featuring consistent data. The first challenge evaluated participants based on the root mean square error (RMSE) obtained and the percentage of correct values within a logS error range. Unfortunately, the authors did not report the methods used by the participants.[27] In contrast, the second challenge showed that regardless participants were open to choosing which approach to use, all submitted responses used an implementation of QSPR or machine learning (ML).[28] Neural networks (NN), multiple linear regression (MLR), and decision trees were the most commonly applied methods in these challenges. Tree-based and MLR models presented the best results. Surprisingly, new state-of-art methods did not yield a significant improvement in predictions compared to the results of the first challenge.[27] The challenges’ findings showed that data quality is more critical for accurate predictions than model selection.[28] The results from the solubility challenges will be discussed in detail in Section 5.
Ideally, solubility models should be accurate and accessible, having clear or minimal instructions on how to use a model. Thus a common idea is to use web servers to provide easier public access. However, maintaining a web server requires an ongoing investment of time and money. There are examples of servers that eventually disappear, even with institutional or government support[29]. For example, eight of the 89 web server tools from the 2020 Nucleic Acid Research special web server issue are already offline[30]111Tested December 30, 2022 after just a few years. Additionally, some tasks may require a long computation time[31]. For instance, tools like RoseTTAFold[32] and ATB[33] can take hours to days to complete a job, resulting in long queues and waiting times.
An alternative approach is to perform the computation directly on the user’s device, removing the need for the server’s maintenance and cost. In this approach, the website is simply a static file that can be hosted on sites like GitHub and be completely archived in the Internet Archive222https://archive.org/. We explored this approach in Ansari and White [34] for bioinformatics. The main drawback is that the application runs directly from the browser on a user’s device (a personal computer or even a cellphone). This would be infeasible for first-principle methods, like those that rely on molecular dynamics. Nevertheless, it is feasible for deep learning models, especially with the increasing integration of deep learning chips and compiler optimizations.
In this work, we developed a front-end application using a JavaScript (JS) implementation of TensorFlow framework[35]. Our application can be used to predict the solubility of small molecules with uncertainty. To calibrate the confidence of the prediction, our model implements a deep ensemble approach[36] which allows reporting model uncertainty when reporting the prediction. Our model runs locally on the user’s device and can be accessed at https://mol.dev/.
2 Related works
Physic-based models have been developed in the past for aqueous solubilization predictions. Those models may become complex, limiting their use to advanced users only.[24] Despite physic-based models being derived from first principles, they are no more accurate than empirical methods.[37] Data-driven models can outperform physics-based models with the benefit of being less time-consuming. Historically, common approaches computed aqueous solubilities based on QSPR[21, 23, 22] and MLR[38, 39] methods.[24]
Huuskonen [38] used a dataset consisting of 1297 organic molecules to develop two models based on MLR and NN. The author reported a good correlation between predicted properties and labels for training () and test () data. Delaney [39] used another approach based on MLR called Estimated SOLubility (ESOL) adjusted on a 2874 small organic molecules dataset. The final model presented a and an average absolute error (AAE) of 0.83. GPsol[40], a Gaussian Process-based model, was trained to predict the aqueous solubilities of electrolytes in addition to non-electrolyte molecules. It used 1664 descriptors computed by Dragon software as input features to train the model on a dataset of 4000 molecules. Depending on the dataset, it presented an RMSE of or . Lusci et al.[41] trained several Undirected graph recurrent neural networks (UG-RNN) architectures using different sets of node feature vectors. The authors report RMSE from 0.90 to 1.41 for different models on the first Solubility Challenge dataset[25]. McDonagh et al.[37] calculated solubilization free energies using first principle theoretical calculations and cheminformatic methods. Their results have shown that cheminformatic methods have better accuracy than theoretical methods. The authors also point to the promising results of using Random Forest models (RMSE of 0.93 on Llinàs first dataset[25]).
Those models use descriptors to represent molecules. Descriptors are a straightforward way to convey physical-chemical information in your input. However, descriptor selection is not an easy task. It requires a good understanding of the problem settings, usually only held by specialists. Some automated methods have been proposed to select descriptors[42]. Nevertheless, computing several descriptors can increase the time needed for inference. Additionally, these descriptors can be valid only for a specific region of the chemical space[43].
@@ 2 Related works
More recently, transformers models have been used to compute the solubility of small molecules. Francoeur et al.[44] developed the SolTranNet, a transformers model trained on AqSolDB[1] solubility data. Notably, this architecture results in an RMSE of only when trained and evaluated on the original ESOL[39] dataset using random split. Nevertheless, it shows an RMSE of when trained using the AqSolDB[1] and evaluated using ESOL. It suggests that the molecules present in ESOL may have low variability, meaning that samples in the test set are similar to samples in the training set. Hence, models trained on the ESOL training set performed excellently when evaluated on ESOL test set. Regression Transformer (RT)[45] is a multipurpose transformer model trained using an infilling mask approach[46]. Their results are comparable to those achieved by the SMILES-BERT[47] and Mol-BERT[48] models. While the RMSE values for SMILES-BERT and Mol-BERT are and , respectively. Whereas RT presented an RMSE of . All three models were fine-tuned using ESOL. MolFormer[49] is an encoder-only transformers model with a modified embedding. It was pre-trained in a large corpus and fine-tuned for numerous downstream tasks. Specifically for the solubility regression fine-tuning, they reported an RMSE of on the ESOL dataset. Noticeably, the same value was reported by Francoeur et al.[44] when they trained their model on ESOL.
Comparing the performance of different models is a complex task, as performance metrics cannot be directly compared across models evaluated on distinct datasets. To address this issue, Panapitiya et al. [50] curated a large and diverse dataset to train models with various architectures and molecular representations. They also compared the performance of these models on datasets from the literature[51, 38, 52, 53, 39, 54, 25, 55, 26, 56, 57, 58]. Although their models achieved an RMSE of on their test set, using descriptors as molecular representations resulted in RMSE values ranging from to when applied to other datasets from the literature. These findings suggest that some datasets used to train models in the literature may be inherently easier to predict, leading to smaller RMSE values. According to their study, the Solubility Challenge datasets by Llinàs et al.[25, 26] were found to be particularly challenging due to their more significant reproducibility error.
@@ 3.1 Dataset
We augmented AqSolDB to 96,625 molecules. Each entry of AqSolDB was used to generate at most ten new unique randomized SMILES strings. Training the model on multiple representations of the same molecule improves its ability to learn the chemical space constraints of the training set, as demonstrated in previous studies [62, 63]. Duplicates were removed.
After shuffling, the augmented dataset was split into 80%/20% for the training and test datasets, respectively. The curated datasets for the solubility challenges[25, 28] were used as withheld validation data to evaluate the model’s ability to predict solubility for unseen compounds. To refer to the validation datasets, we labeled the first solubility challenge dataset as "solubility challenge 1" and the two sets from the second solubility challenge as "solubility challenge 2_1" and "solubility challenge 2_2", respectively. Molecules present in these three datasets were not found in train and test datasets.
@@ 3.2 Model architecture
Our model uses a deep ensemble approach as described by Lakshminarayanan et al. [36]. Given a model which outputs two values (mean and variance ), a deep ensemble creates an ensemble of models that can estimate prediction uncertainty. Those two numbers characterize a normal distribution , where indexes the model in the ensemble.
The uncertainty of a model can be divided into two sources: aleatoric uncertainty (AU) and epistemic uncertainty (EU).[64, 65] EU quantifies the uncertainty among the models of the ensemble. It shows how much the elements of the ensemble disagree about a prediction. EU is also known as “model uncertainty”. AU, also called “data uncertainty”, quantifies intrinsic uncertainty inherent in data observations.[66]
For a given data point , the estimates for the ensemble predictions are computed as follows:
@@
As the base model for deep ensembling, we built a deep neural network (DNN) using Keras[67] framework and TensorFlow[68] back-end. Given its capabilities to capture long-range sequence correlations, we employed a bidirectional recurrent neural network (RNN) layers in our model. Figure 1 illustrates the model architecture.
The input Simplified molecular-input line-entry system (SMILES)[69] or Self-referencing embedded strings (SELFIES)[70] molecule representation is converted to a list of SELFIES tokens according to a pre-defined vocabulary. The vocabulary was created based on the training data, generating 273 available tokens. SELFIES tokens are the input for the DNN.
The network can be divided into three sections: () Embedding, () bi-RNN, and () fully connected NN.
@@
As the inset of Figure 1 shows, the first step is to input the SELFIES tokens into an embedding layer. The embedding layer allows us to convert a list of discrete tokens into a fixed-length vector space. Working on a continuous vector space has two main advantages: it uses a more compact representation, and semantically similar symbols can be described close to each other in vector space. Our embedding layer has an input dimension of 273 (vocabulary size) and an output dimension of 64.
@@
After generating the embedding vector, the next step is to feed it into the bidirectional RNN layer. The effects of using Gated Recurrent Unit (GRU) or Long Short-Term Memory (LSTM)[71] layers as the RNN layers were investigated. It will be shown that LSTM performed better for this application (refer to Section 4.1). The bi-LSTM layer consists of a double-stacked LSTM layer of 64 units each. These layers use three gates (input, forget, and output gates) to filter information (See Ref 72 for details). Using bi-RNN was motivated based on our previous work[34] in which LSTM helped improve the model’s performance for predicting peptide properties using its sequences.
The output from the bi-LSTM stack undergoes normalization via Layer Normalization[73]. In DNN models, gradient values concerning weights from one layer heavily rely on outputs from the preceding layer, an issue referred to as "covariate shift.". Some authors argue that normalization schemes improve the model by reducing covariate shifts during forward normalization.[74, 75] Conversely, others argue that improvements are based on the derivatives of the mean and variance by effectively re-scaling gradients.[76, 77] The absence of a comprehensive theoretical grasp of normalization effects hinders the evolution of novel regularization schemes.[78] Furthermore, normalization serves to re-center layer values around 0, where the non-linearity of most activation functions is more intense. Despite the limited understanding, Layer Normalization is employed due to its demonstrated effectiveness.[77]
@@
In the final section of the DNN, normalized data is processed through three dense layers containing 32, 16, and 1 units, respectively. However, the output of the dense layer with 16 units is fed separately into two separate dense layers with 1 unit each. One layer employs a linear activation function, while the other utilizes a softplus activation function to ensure a positive value, producing and , respectively.
To avoid divergences caused by a division by zero or convergence to large uncertainties, the variance value is constrained between the minimum and the maximum values of 10 and , respectively. The outputs of these two last layers are stacked to generate the model output. Therefore, the model produces a tuple with the mean and the variance. Negative log-likelihood loss was used to train the model. It is defined as the probability of observing the label given the input :
@@
During the training phase, dropout layers with 0.35 dropout rate were incorporated after the embedding and each dense layer to mitigate over-fitting.[79] Figure 1 illustrates the model applied for inference, which omits the dropout layers. Models were trained using the Adam[80] optimizer with a fixed learning rate of 0.0001 and default values for and (0.9 and 0.999, respectively).
@@
Our model employs adversarial training, following the approach proposed by Lakshminarayanan et al. [36]. However, our input is a tokenized version of the SELFIES representation. Hence, it is a discrete sequence. To apply adversarial training, we generate adversarial examples by modifying the embedded representation of the input data. Each iteration in the training phase consists of first computing the loss using Equation 3 and a second step with a new input to smooth the model’s prediction:
@@
Details of the model are available as model cards[81] at http://mol.dev/. These cards provide information concerning performance, limitations, training data, ethical considerations, and caveats of the model.
@@ 4 Results
In order to evaluate the performance of our model using deep ensembles, two baseline models were created: () an XGBoost Random Forest (RF) model using the 17 descriptors available on AqSolDB plus 1809 molecular descriptors calculated by PaDELPy, a python wrapper for the PaDEL-Descriptor[82] software, and () a model with the same architecture used on our deep ensemble using RMSE as the loss function and no ensemble (referred to as DNN). In addition, we evaluate the influence of () the bi-RNN layer (either GRU or LSTM), () using an augmented dataset to train, () the adversarial training, and () the ensemble size.
@@ 4.1 Gated layer
We trained models using GRU or LSTM in bidirectional layers selecting the model using a deep ensemble with four individual models. Metrics can be found in Table 1; for an explanation of the naming syntax used in this work, refer to Table 1 caption. Using LSTM resulted in a decrease in RMSE and MAE and an increase in the correlation coefficient, indicating better performance. For Solubility Challenges 1, 2_1, and 2_2, the kde4 model yielded RMSE values of 1.329, 1.354, and 1.626, respectively, while the kde4 model achieved 1.049, 1.054, and 1.340, respectively. This trend was also observed for the models trained without data augmentation, but in a smaller proportion (See Table 1). Considering that LSTM performs better regarding this model and data, we will consider only bi-LSTM layers for further discussion.
@@ 4.2 Data augmentation
Our model is not intrinsically invariant with respect to the selfies representation input. For instance, both “C(C(C1C(=C(C(=O)O1)O)O)O)O” and “O=C1OC(C(O)CO)C(O)=C1O” are valid SMILES representations for the ascorbic acid (See Figure 1) that will be encoded for different SELFIES. Hence, the model should learn to be invariant concerning changes in the string representation during training. It can be achieved by training the model using different representations with the same label. Therefore, the model can learn relations in the chemical space instead of correlating the label with a specific representation. With this aim, we evaluated the effects of augmenting the dataset by generating new randomized SMILES representations for each sample.[62]
Among the performance tests, augmenting the dataset had the most significant impact on RMSE. It could be seen improvements of in the RMSE when evaluating on challenge datasets 1 and 2_1, and an improvement of on 2_2 (See Table 1). Concerning the first two datasets, augmenting data improved every model used in this study. However, surprisingly, data augmentation led to a deprecation of the DNN model on the solubility challenge 2_2 dataset. This behavior was not further investigated.

### v4

# Predicting small molecules solubility on endpoint devices using deep ensemble neural networks

@@ 1 Introduction
Solubility prediction is essential, driving the development of various methods, from physics-based approaches — including first principles[10, 11], semi-empirical equations[12, 13, 14], molecular dynamics (MD)[15, 16, 17, 18], and quantum computations[19] — to empirical methods like quantitative structure-property relationship (QSPR)[20, 21, 22, 23] and multiple linear regression (MLR)[24, 25]. Despite their sophistication, physics-based models often present complexity that limits accessibility to advanced users[26] and do not guarantee higher accuracy than empirical methods[27]. Data-driven models emerge as efficient alternatives, capable of outperforming physics-based models[26]. However, achieving accurate and reliable solubility predictions remains a significant challenge[26, 28].
To address the persistent issues of systematic bias and non-reproducibility in aqueous solubility datasets, Llinàs et al.[29, 30] introduced two solubility challenges featuring consistent data. The first challenge red participants based on the root mean square error (RMSE) and the accuracy within a logS error range[31]. The second challenge revealed that despite the freedom in method selection, all entries relied on QSPR or machine learning (ML) techniques[32], yet did not achieve a notable improvement over the first challenge[31]. These challenges highlighted the importance of data quality over model selection for accurate solubility predictions.[32] Sorkun et al. [28] further emphasized this by demonstrating how data quality assessments on subsets of the AqSolDB[1] significantly impacted model performance[28].
McDonagh et al. [27] demonstrated that cheminformatic methods surpass first principle theoretical calculations in calculating solubilization free energies, highlighting the superior accuracy of cheminformatics and the efficacy of Random Forest models, evidenced by an RMSE of 0.93 using Llinàs’ first dataset[29]. Data-driven approaches, particularly feature-based models, have contributed to accurate aqueous solubility prediction. Delaney [25] used MLR to develop a model called Estimated SOLubility (ESOL) adjusted on a 2874 small organic molecules dataset with an average absolute error (AAE) of 0.83. Comparable performance has been achieved using various methods including MLR[24], Gaussian processes[33], undirected graph recurrent neural networks (UG-RNN)[34], deep neural networks (DNN)[35], and random forests (RF)[36, 37].
@@
Recently, transformers[38] models have been applied to compute solubility of small molecules.[39, 40, 41, 42, 43]. Francoeur and Koes [41] developed the SolTranNet, a transformers model trained on AqSolDB[1] solubility data. Notably, this architecture results in an RMSE of only when trained and evaluated on the original ESOL[25] dataset using random split. Nevertheless, it shows an RMSE of when trained using the AqSolDB[1] and evaluated using ESOL. It suggests that the molecules present in ESOL may have low variability, meaning that samples in the test set are similar to samples in the training set. Hence, models trained on the ESOL training set performed excellently when evaluated on the ESOL test set.
Solubility models should ideally combine accuracy with ease of access. Thus, a common idea is to use web servers to provide easier public access. However, web servers demand continuous financial and time investments for maintenance, leading to the eventual disappearance of some, despite having institutional or government backing[44]. For instance, eight out of 89 web server tools featured in the 2020 Nucleic Acids Research special web server issue were offline by the end of 2022[45]. Moreover, computational demands can be significant, with tools like RoseTTAFold[46] and ATB[47] requiring hours to days for job completion, thus creating potential delays due to long queues and wait times[48].
An alternative approach is to perform the computation directly on the user’s device, removing the need for the server’s maintenance and cost. This method allows hosting the website as a static file on platforms such as GitHub, with potential archiving on the Internet Archive111https://archive.org/. We explored this approach in Ansari and White [49] for bioinformatics. Our web application implements a deep ensemble[50] recurrent neural network (RNN) capable of extracting data directly from molecular string representations, such as SMILES[51] or SELFIES[52], which can be easily quickly accessed.[53, 54]
The primary difficulty lies in the application’s dependence on the device’s capabilities, which is crucial for smartphones with limited resources. Balancing performance in low-resource settings, the use of transformer models[38] becomes impractical due to their large size, incompatible with smartphone memory and prolonged inference times. Additionally, our model implements a deep ensemble to calibrate uncertainties, making the application of transformers even more unfeasible. In contrast, using descriptors is an easy way to convey physical information to the model and, consequently, enables smaller models. However, descriptor computation is time-intensive. In our tests, using PaDEL to compute descriptors for all molecules in AqSolDB took roughly 20 hours. Furthermore, feature-based model development requires specialized knowledge for feature selection[55], and is limited by the regions of the chemical space these descriptors cover[56]. Even application usage may need specialized data, as Kurotani et al. [37] illustrate. RNNs present an alternative for property extraction directly from string representations while allowing for adaptable computational resource management.
In this work, we developed a front-end application using a JavaScript (JS) implementation of TensorFlow framework[57]. Our application can be used to predict the solubility of small molecules with uncertainty. To calibrate the confidence of the prediction, our model implements a deep ensemble approach[50] which allows reporting model uncertainty when reporting the prediction. Our solution implements a deep ensemble of RNN models specially designed to achieve satisfactory performance while being able to run in an environment without strong computational resources. This application runs locally on the user’s device and can be accessed at https://mol.dev/. Mol.dev does not save data input for predictions in any way.
@@ 2.1 Dataset
We augmented AqSolDB to 96,625 molecules using SMILES randomization[61, 62]. Each entry of AqSolDB was used to generate at most ten new unique randomized SMILES strings. Training the model on multiple representations of the same molecule improves its ability to learn the chemical space constraints of the training set, as demonstrated in previous studies [61, 62]. Duplicates were removed.
After shuffling, the augmented dataset was split into 80%/20% for the training and test datasets, respectively. The curated datasets for the solubility challenges[29, 32] were used as withheld validation data to evaluate the model’s ability to predict solubility for unseen compounds. To refer to the validation datasets, we labeled the first solubility challenge dataset as "solubility challenge 1" and the two sets from the second solubility challenge as "solubility challenge 2_1" and "solubility challenge 2_2", respectively. Molecules in these three datasets were not found in train and test datasets.
@@ 2.2 Model architecture
Our model uses a deep ensemble approach as described by Lakshminarayanan et al. [50]. This technique was selected due to its ability to estimate prediction uncertainty, thus enhancing the predictive capability of our model. The uncertainty of a model can be divided into two sources: aleatoric uncertainty (AU) and epistemic uncertainty (EU).[63, 64] These uncertainties quantify the intrinsic uncertainty inherent in data observations and the disagreement among model estimations, respectively.[65]
Given a model that outputs two values – and – that characterize a normal distribution , a deep ensemble creates an ensemble of models that can estimate prediction uncertainty. For a given data point , the estimates for the ensemble predictions are computed as follows:
@@
We used a deep neural network (DNN) implemented using Keras[66] and TensorFlow[67] to build the deep ensemble. Our DNN model uses Self-referencing embedded strings (SELFIES)[52] tokens as input. A pre-defined vocabulary was created by analyzing all training data. Each unique SELFIES group was assigned to a corresponding integer, yielding 273 distinct tokens. Simplified molecular-input line-entry system (SMILES)[51] or SELFIES[52] molecule representations are converted to tokens based on the pre-defined vocabulary. Figure 1 illustrates the model architecture. The network can be divided into three sections: () Embedding, () bi-RNN, and () fully connected NN.
@@
The embedding layer converts a list of discrete tokens into a fixed-length vector space. Working on a continuous vector space has two main advantages: it uses a more compact representation, and semantically similar symbols can be described closely in vector space. Our embedding layer has an input dimension of 273 (vocabulary size) and an output dimension of 64.
@@
Following the embedding layer, the data are fed into the bidirectional Recurrent Neural Network (RNN) layer. We used two RNN layers, each containing 64 units. The effects of using Gated Recurrent Unit (GRU) or Long Short-Term Memory (LSTM)[68] layers as the RNN layers were investigated (refer to Section 3.1). Using bi-RNN was motivated based on our previous work[49] in which LSTM helped improve the model’s performance for predicting peptide properties using its sequences. More details regarding RNN, LSTM, and GRU layers can be found in Ref. 69.
The output from the bi-LSTM stack undergoes normalization via Layer Normalization[70]. There is no agreement on why Layer Normalization improves the model’s performance.[71, 72, 73, 74] The absence of a comprehensive theoretical understanding of normalization effects hinders the evolution of novel regularization schemes.[75] Despite the limited understanding, Layer Normalization is employed due to its demonstrated effectiveness.[74]
@@
After normalization, data is processed through three dense layers containing 32, 16, and 1 units, respectively. The 16-unit layer’s output goes to two different 1-unit layers. One layer uses a linear function and the other uses a softplus function, producing and , respectively.
Negative log-likelihood loss was used to train the model. It is defined as the probability of observing the label given the input :
@@
During the training phase, dropout layers with 0.35 dropout rate were incorporated after the embedding and each dense layer to mitigate over-fitting.[76] Models were trained using the Adam[77] optimizer with a fixed learning rate of 0.0001 and default values for and (0.9 and 0.999, respectively).
@@
Our model employs adversarial training, following the approach proposed by Lakshminarayanan et al. [50] to improve the robustness of our model predictions. Because the input for our model is a discrete sequence, we generate adversarial examples by modifying the embedded representation of the input data. Each iteration in the training phase consists of first computing the loss using Equation 3 and a second step with a new input to smooth the model’s prediction:
@@
Details of the model performance, limitations, training data, ethical considerations, and caveats are available as a model card[78] at http://mol.dev/.
@@ 3 Results
In order to evaluate the performance of our model using deep ensembles, two baseline models were created: () an XGBoost Random Forest (RF) model using the 17 descriptors available on AqSolDB plus 1809 molecular descriptors calculated by PaDELPy, a python wrapper for the PaDEL-Descriptor[79] software, and () a model with the same architecture used on our deep ensemble using RMSE as the loss function and no ensemble (referred to as DNN). RFs are the state-of-the-art (SOTA) of solubility prediction. We used this baseline as a comparison to prove that our model is able to achieve SOTA performance using only molecular string representations. In addition, we evaluate the effects of () the bi-RNN layer (either GRU or LSTM), () using an augmented dataset to train, () the adversarial training, and () the ensemble size in the model’s performance. Table 1 shows the performance of each one of our trained models.
@@ 3.1 Gated layer
We trained models with four elements in the deep ensemble using GRU or LSTM. Metrics can be found in Table 1; for an explanation of the naming syntax used in this work, refer to Table 1 caption. Using LSTM resulted in a decrease in RMSE and MAE and an increase in the correlation coefficient, indicating better performance. For Solubility Challenges 1, 2_1, and 2_2, the kde4 model yielded RMSE values of 1.329, 1.354, and 1.626, respectively, while the kde4 model achieved 1.273, 1.137, and 1.511, respectively. This trend was also observed for the models trained without data augmentation (See Table 1). Considering that LSTM performs better regarding this model and data, we will consider only bi-LSTM layers for further discussion. Those results are in accordance with our previous work[49] in which using LSTM helped improve the model’s performance.
@@ 3.2 Data augmentation
Our model is not intrinsically invariant with respect to the SELFIES representation input. For instance, both “C(C(C1C(=C(C(=O)O1)O)O)O)O” and “O=C1OC(C(O)CO)C(O)=C1O” are valid SMILES representations for the ascorbic acid (See Figure 1) that will be encoded for different SELFIES tokens. Hence, the model should learn to be invariant concerning changes in the string representation during training. It can be achieved by augmenting the dataset with SMILES randomization and training the model using different representations with the same label. Therefore, the model can learn relations in the chemical space instead of correlating the label with a specific representation.[61] With this aim, we evaluated the effects of augmenting the dataset by generating new randomized SMILES representations for each sample.
Augmenting the dataset had a significant impact on the metrics. It could be seen improvements of in the RMSE when evaluating on challenge datasets 1 and 2_1, and a gain of on 2_2 (See Table 1). Concerning the first two datasets, augmenting data improved every model used in this study. However, surprisingly, data augmentation led to a deprecation of the DNN model on the solubility challenge 2_2 dataset. This behavior was not further investigated.